# Krylov Complexity and Dynamical Phase Transition in the quenched LMG model

### v1

@@ IV DPT-I in the LMG model
where $S_{\alpha}=\sum_{i=1}^{N}\sigma^{\alpha}_{i}/2$ ($\alpha=x,y,z$) are collective spin operators, with $\sigma^{\alpha}_{i}$ denoting the $\alpha$ Pauli matrix acting on the $i$-th site of a chain of $N=2j$ sites and total angular momentum $j$. This Hamiltonian is a particular case of the well-known Lipkin-Meshkov-Glick model [36]. The constant $J$ denotes the ferromagnetic coupling between the spins in the $z$ direction while $h$ represents the strength of magnetic field applied along the $x$-axis. This Hamiltonian describes an ensemble of $N$ spin-$1/2$ systems subject to all-to-all pairwise interactions in the presence of a magnetic field.
Considering a quench $h_{0}=0\rightarrow h_{f}>0$, we show in Fig. 1(a) the bifurcation diagram of Hamiltonian (14) for the time-averaged magnetization (order parameter for this model)
@@
Figure 1: Magnetization as an order parameter. Panel (a) shows the time-averaged magnetization $\overline{S_{z}}$ as a function of quench $h_{0}=0\rightarrow h_{f}=h$. The initial state is $|\psi_{0}\rangle=|\downarrow\rangle_{z}$. In the thermodynamic limit, $\overline{S_{z}}$ signals the dynamical phase transition at the critical point $h=1/2$, thus identifying two dynamical critical phases. Panels (b) and (c) show the magnetization $S_{z}(t)$ as a function of time for a quench in the ferromagnetic phase ($h=0.3$) and a quench crossing the critical point to the paramagnetic phase ($h_{f}=0.8$), respectively. In both phases, the thermodynamic limit prevents the magnetization from revivals.
@@
An essential aspect of DPT-I in this context is the symmetry of the initial state. This critical phenomenon is only observed when the initial state of the dynamics exhibits broken symmetry [48]. The LMG model feature spin-flip symmetry [62], represented by the operator $S=e^{i\pi(S_{x}-j)}$. This symmetry dictates that $H$ does not couple standard eigenbasis $|j,m_{z}\rangle$ states (eigenvectors of $S^{2}$ and $S_{z}$) with even and odd $m_{z}$. Under these conditions, a quench within the same dynamical phase produce oscillations in the system around a broken-symmetry effective state, while a quench crossing the dynamical critical point produces oscillations around a symmetric effective state. These oscillations are characterized by long-lived steady states and they are usually linked to prethermalization [63].
Figures 1(b) and 1(c) illustrate instances of magnetization behavior over time, depicting a quench within the same phase and crossing the dynamical critical point, respectively. It is evident that the observed revivals in these figures arise from finite-size effects within the system. As the system size $N$ increases, these revivals disappear. In the following section, we demonstrate that the time-averaged Krylov complexity can also serve as an order parameter for this system.
@@ V Results
We start by discussing the behaviour of the spread complexity and its long-time average. In order to understand why Krylov complexity can be taken as an order parameter for the DPT-I in the LMG model, we proceed with the analysis of two other quantities, the inverse participation ratio and the Shannon entropy, considering two distinct basis, the energy and the Krylov basis.
@@ V.1 Krylov Complexity of quantum states in the LMG model
Figure 2: Krylov complexity as an order parameter. Panel (a) shows the normalized time-averaged Krylov complexity $\overline{C}$ as a function of $h$ considering $h_{0}=0$ and initial state $|\psi_{0}\rangle=|\downarrow\rangle_{z}$. $\overline{C}$ exhibits exactly the same transition as $\overline{S_{z}}$ at the critical point $h_{c}=1/2$. Panels (b) and (c) show the Krylov complexity $C_{\mathcal{K}}(t)$ as a function of time for a quenches in the ferromagnetic phase ($h_{f}=0.3$) and across the critical point, to the paramagnetic phase ($h_{f}=0.8$). In both phases, the thermodynamic limit also prevents the complexity from revivals.
@@
Using the same setup described in the last section, Fig. 2(a) shows the normalized time-averaged complexity as function of the quench intensity $h$. Remarkably, $\overline{C}$ exhibits the same qualitative behaviour as the dynamical order parameter $\overline{S_{z}}$, with the only difference being the numerical values assumed in both dynamical phases.
Similar to the magnetization $S_{z}(t)$, the Krylov complexity exhibits distinct behavior in each dynamical phase over time. This characteristic is depicted in panels (b) and (c) of Fig. 2. Extensive numerical calculations suggest that, in the thermodynamic limit, and for $h_{0}=0$, the complexity oscillates around a finite value, consistently remaining below unity for a quench within the ferromagnetic phase. Conversely, it oscillates around the unit value for a quench into the paramagnetic phase, thus crossing the critical point $h_{c}=1/2$. Based on these findings, we can argue that we use the time-averaged Krylov complexity to indicate a DPT-I.
@@
Given that DPT-I is associated with a dynamic symmetry break, it is reasonable to anticipate that the sensitivity of the Krylov basis extends to the symmetry of the model. Therefore, a deeper relationship between the Krylov basis and the energy one should exist. This aspect will be investigated in the concluding part of this section.
@@ V.2 Inverse Participation Ratio
In order to analyze the relation between the energy basis and the Krylov basis, we first consider the inverse participation ratio (IPR) in each one of these bases. The IPR is given by
$IPR(t)=\sum_{k}|\langle k|\psi(t)\rangle|^{4}=\sum_{k}p_{k}^{2}(t),$ (17)
for some basis whose elements we generically denote as $\{|k\rangle\}$. As its name suggests, the IPR measures how many states of the chosen basis effectively participate in the course of the time evolution of the system.
Considering the same protocol as before, we compute the IPR for two distinct bases: the pre-quench energy eigenbasis (the eigenvectors of $H_{0}$) and the Krylov basis. Two instances of the results are shown in Fig. 3. Interestingly, the IPR computed in both bases are identical. However, this coincidence only happens when we start with $h_{0}=0$, when the initial energy spectra is doubly degenerate. Extensive numerical analysis shows that they are different otherwise. Another property of IPR that we observed numerically is the independence with respect to which ground-state is taken as initial state, thus giving the same result also for the second ground-state $|\psi_{0}\rangle=|\uparrow\rangle_{z}$.
Figure 3: Inverse Participation Ratio. Panel (a) shows IPR for a quench $h=0.3$, while panel (b) shows the same quantity, but for a quench crossing the critical point ($h=0.8$). We take $N=200$ in both panels. Exactly the same behaviour is observed if the initial state is taken as the second ground-state, $|\uparrow\rangle_{z}$.
The reason for this match will be analytically explored latter in this article. However, before presenting this analysis, let us consider another interesting quantity, the Shannon entropy.
@@ V.3 $\mathcal{K}$-Entropy - Shannon entropy in the Krylov basis
is a measure of the uncertainty of the system —or the classical information contained in the system. We are here interested in the entropy associated with both bases discussed in the previous section: the Krylov basis, for which $p_{n}(t)=|\langle K_{n}|\psi(t)\rangle|^{2}$ and the initial energy basis, whose probabilities are $p_{n}(t)=|\langle E^{0}_{n}|\psi(t)\rangle|^{2}$. The Shannon entropy in the Krylov basis was introduce in Ref. [12] and is known also as the $\mathcal{K}$-entropy. It is a measure of complexity of the dynamics. The authors of Ref. [31] argued that the complexity defined as the exponential of this quantity measures the minimum Hilbert space dimension required to store the probability distribution $p_{n}(t)$.
By comparing the Shannon entropy in both basis we again observe a perfect match. Also, such feature only occurs when we start at $h_{0}=0$. An example is shown in Fig. 4 for the same parameters used in Fig. 3.
@@
We note that the probabilities $p_{n}(t)=|\langle E^{0}_{n}|\psi(t)\rangle|^{2}$ are the populations of the density matrix in the energy eigenbasis and, thus, the entropy associated with such distribution is the diagonal entropy, proposed as the thermodynamic entropy for closed quantum systems [64]. Since $p_{n}(t)=|\langle K_{n}|\psi(t)\rangle|^{2}$ can be seen as the probability density associated to site $n$ in the semi-infinite Krylov lattice, naturally the corresponding Shannon entropy $\mathcal{E}_{\mathcal{K}}(t)$ can be interpreted as the uncertainty in the spreading of the initial state through the Krylov subspace.
The numerical results described above indicates a deeper match, beyond average spread. The goal of the next subsection is to investigate why this happens.
V.4 Derivation of the Krylov basis
@@
$H_{f}|j,m_{z}\rangle=c_{0}|j,m_{z}\rangle+c_{+}|j,m_{z}+1\rangle+c_{-}|j,m_{z}-1\rangle$ (22)
@@
The above discussion proves that the Krylov complexity must have the same time behaviour as the magnetization. We thus conclude that its time average is an order parameter for this model.
Finally, we note that when $|\psi_{0}\rangle=|j,m_{z}\rangle$ for any $m_{z}$, $\dim{(\mathcal{K})}=\dim{(\mathcal{H})}$, i.e., the dimension of the Krylov subspace is equal to the dimension of the Hilbert space.These initial states explore the whole Hilbert space, showing some kind of ergodicity. We emphasize that this is not an feature of all quantum systems.
@@ VI Conclusion
Understanding the complex nature of the temporal evolution of many-body quantum systems holds fundamental significance across various research fields. This study contributes to this area by examining the Krylov complexity (spread complexity) and its connection to the dynamical phase transition within the LMG model. In essence, we show that the time-averaged Krylov complexity serves as an order parameter in this context. This conclusion stems from a numerical investigation encompassing not only the Krylov complexity but also the inverse participation ratio and Shannon entropy within the Krylov basis. These analysis suggest a deep relation between the Krylov basis and the energy eigenbasis. Subsequently, an analytical study establishes that the time-averaged Krylov complexity effectively signals the dynamical phase transition in this model by means of the equivalence between the Krylov and the energy bases.
It is important to note that a thorough numerical investigation reveals that this connection is valid exclusively in instances where the symmetry of the model is broken. Our analytical analysis further substantiates this conclusion. This prompts questions about the how Krylov complexity behaves under general changes of symmetry. Addressing this question will certain deepen our comprehension of the dynamics inherent in many-body systems.
Our findings also provide insights into the thermodynamics of critical systems of this nature. While the thermodynamics of equilibrium phase transitions is well-established, the same cannot be said for its dynamic counterpart. Our results indicate that, in the case of the LMG model, the $\mathcal{K}$-entropy serves as the quantum thermodynamic entropy. This assertion is grounded in the definition of diagonal entropy [64] and the gauge theory of quantum thermodynamics outlined in Ref.[65]. Furthermore, this observation aligns with earlier investigations into the thermodynamics of dynamical quantum phase transitions, where a thermodynamic entropy in quantum phase space not only signals the transition but also, on average, exhibits a monotonic increase over time [66]. Consequently, our results suggest a profound connection between dynamical critical behavior and the process of thermalization.
@@
Whether the time-averaged Krylov complexity can function as an order parameter for DPT-I in systems beyond the LMG model remains an open question. Additionally, delving into the broader applicability of the connection between Krylov and energy bases is crucial. Addressing these questions promises to provide valuable insights into the underlying nature of this dynamic critical behavior.

### v2

@@ IV DPT-I in the LMG model
where $S_{\alpha}=\sum_{i=1}^{N}\sigma^{\alpha}_{i}/2$ ($\alpha=x,y,z$) are collective spin operators, with $\sigma^{\alpha}_{i}$ denoting the $\alpha$ Pauli matrix acting on the $i$-th site of a chain of $N=2j$ sites and total angular momentum $j$. This Hamiltonian is a particular case of the well-known Lipkin-Meshkov-Glick model [40]. The constant $J$ denotes the ferromagnetic coupling between the spins in the $z$ direction while $h$ represents the strength of the magnetic field applied along the $x$-axis. This Hamiltonian describes an ensemble of $N$ spin-$1/2$ systems subject to all-to-all pairwise interactions in the presence of a magnetic field.
Considering a quench $h_{0}\rightarrow h_{f}>0$, we show in Fig. 1(a) the bifurcation diagram of Hamiltonian (14) for the time-averaged magnetization (order parameter for this model)
@@
Figure 1: Magnetization as an order parameter. Panel (a) shows the time-averaged magnetization $\overline{S_{z}}$ as a function of quench $h_{0}=0\rightarrow h_{f}=h$. The initial state is $|\psi_{0}\rangle=|\downarrow\rangle_{z}$. In the thermodynamic limit, $\overline{S_{z}}$ signals the dynamical phase transition at the critical point $h=1/2$, thus identifying two dynamical critical phases. Panels (b) and (c) show the magnetization $S_{z}(t)$ as a function of time for a quench in the ferromagnetic phase ($h=0.3$) and a quench crossing the critical point to the paramagnetic phase ($h_{f}=0.8$), respectively. In both phases, revivals are suppressed in the thermodynamic limit.
@@
An essential aspect of DPT-I in this context is the symmetry of the initial state. This critical phenomenon is only observed when the initial state of the dynamics exhibits broken symmetry [53]. The LMG model feature spin-flip symmetry [69], represented by the operator $\hat{\Pi}=e^{i\pi(\hat{S}_{x}-j)}$. This symmetry dictates that $H$ does not couple standard eigenbasis $|j,m_{z}\rangle$ states (eigenvectors of $S^{2}$ and $S_{z}$) with even and odd $m_{z}$. Under these conditions, a quench within the same dynamical phase produces oscillations in the system around a broken-symmetry effective state, while a quench crossing the dynamical critical point produces oscillations around a symmetric effective state. These oscillations are characterized by long-lived steady states, and they are usually linked to prethermalization [70].
Figures 1(b) and 1(c) illustrate instances of magnetization behavior over time, depicting a quench within the same phase and crossing the dynamical critical point, respectively. The observed revivals in these figures arise from finite-size effects within the system. As the system size $N$ increases, these revivals disappear. In the following section, we demonstrate that the time-averaged Krylov complexity can also serve as an order parameter for this system.
@@ V Results
We start by discussing the spread complexity’s behavior and its long-time average. To understand why Krylov complexity can be taken as an order parameter for the DPT-I in the LMG model, we proceed with the analysis of two other quantities, the inverse participation ratio, and the Shannon entropy, considering the energy and the Krylov basis.
@@ V.1 Krylov Complexity of quantum states in the LMG model
Figure 2: Krylov complexity as an order parameter. Panel (a) shows the normalized time-averaged Krylov complexity $\overline{C}/j$ as a function of $h_{f}$ for several values of $h_{0}$ and $N=200$. $\overline{C}$ exhibits exactly the same transition as $\overline{S_{z}}$ at the dynamical critical point given by $h_{c}=(h_{0}+J)/2$ and represented by the vertical lines. Panel (b) shows the normalized time-averaged Krylov complexity $\overline{C}/j$ for $h_{0}=0.1$ and several values of $N$ as a function of $h_{f}$. As $N$ increases, the non-analytical point of $\overline{C}$ approaches the dynamical critical point $h_{c}=(h_{0}+J)/2=0.55$ (black vertical line), indicating that $\overline{C}$ exhibits the expected behavior of an order parameter in the thermodynamic limit ($N\to\infty$), while the inset shows its normalized derivative w.r.t. $h_{f}$ reaching higher values at $h_{c}$ as $N$ increases. Panels from (c) to (f) show examples of $C_{\mathcal{K}}(t)$ as a function of time for different quenches within the same phase ((c),(d)) and crossing $h_{c}$ ((e),(f)). We see that $C_{\mathcal{K}}(t)$ exhibits the same pattern of oscillations as $S_{z}(t)$. In all panels, we used $|\psi_{0}\rangle=|\downarrow\rangle_{z}$, and in panels (a) and (b), we used $T=150$.
@@
Using the setup described in the last section, Fig. 2(a) shows the normalized time-averaged complexity as a function of the quenching intensity $h_{0}<h_{f}\leq 1$ for several values of $h_{0}$. Remarkably, $\overline{C}$ exhibits the same qualitative behavior as the dynamical order parameter $\overline{S_{z}}$, with the only difference being the numerical values assumed in both dynamical phases.
Moreover, we note that $\overline{C}/j$ exhibits the same dynamical critical point as $\overline{S_{z}}$, whose dependence with $h_{0}$ is $h_{c}=(h_{0}+J)/2$. This is ensured by the fact that as $N$ increases, the non-analytical point of $\overline{C}$ approaches the exact value $h_{c}$, as can be seen from Fig. 2(b) for an example with $h_{0}=0.1$, which implies that as $N$ increases the derivative of $\overline{C}$ with respect to $h_{f}$ reaches even higher values and shall diverge in the thermodynamic limit $N\to\infty$ (see the inset at Fig. 2(b)). This fact was also verified for other values of $h_{0}$ and holds true.
As with the magnetization $S_{z}(t)$, the Krylov complexity exhibits distinct behavior in each dynamical phase over time. This characteristic is depicted in the panels (c)-(f) of Fig. 2. Extensive numerical calculations suggest that, in the thermodynamic limit, and for $h_{0}=0$, the complexity oscillates around a finite value, consistently remaining below unity for a quench within the ferromagnetic phase. Conversely, it oscillates and later stagnates on the unit value for a quench into the paramagnetic phase. As we increase $h_{0}$, we observe the same pattern for quenches within the same phase, however it stagnates on progressively smaller values than unity, hence making its time-average, $\overline{C}$, display smaller plateaus as can be seen from Fig. 2(a) for $h_{f}>h_{c}$. Despite a smaller plateau, $\overline{C}$ seems to reach a constant and maximum value for $h_{f}>h_{c}$, which suggests that DPT-I causes the maximum spreading regime of the dynamics. Based on these findings, we can argue that we use the time-averaged Krylov complexity to indicate a DPT-I.
@@
Given that DPT-I is associated with a dynamical symmetry breaking, it is reasonable to anticipate that the sensitivity of the Krylov basis extends to the symmetry of the model. Therefore, a deeper relationship between the Krylov basis and the energy one should exist. This aspect will be investigated in the concluding part of this section.
Figure 3: Inverse Participation Ratio. Panel (a) shows IPR for a quench $h_{f}=0.3$, while panel (b) shows the same quantity but for a quench crossing the critical point ($h_{f}=0.8$). We take $N=200$ in both panels. Exactly the same behavior is observed if the initial state is taken as the second ground-state, $|\uparrow\rangle_{z}$.
@@ V.2 Inverse Participation Ratio
To analyze the relation between the energy basis and the Krylov basis, we first consider the inverse participation ratio (IPR) in each one of these bases. The IPR is given by
$\text{IPR}(t)=\sum_{k}|\langle k|\psi(t)\rangle|^{4}=\sum_{k}p_{k}^{2}(t),$ (17)
for some basis whose elements we generically denote as $\{|k\rangle\}$. As its name suggests, the IPR measures how many states of the chosen basis effectively participate in the course of the time evolution of the system. Using the analogy with the Krylov chain, it means that the IPR should measure the level of localization of $|\psi_{t}\rangle$ in this chain for a non-trivial time evolution since at $t=0$, $|\psi_{0}\rangle$ is sharply localized in the first site of the Krylov chain as demands the first step of the Lanczos algorithm.
Considering the same protocol as before, we compute the IPR for two distinct bases: the pre-quench energy eigenbasis (the eigenvectors of $H_{0}$) and the Krylov basis. Two instances of the results are shown in Fig. 3. Interestingly, the IPR computed in both bases are identical. However, this coincidence only happens when we start with $h_{0}=0$, when the initial energy spectra is doubly degenerate. Extensive numerical analysis shows that they are different otherwise, see appendix A for some examples of this mismatch. Another property of IPR that we observed numerically is the independence with respect to which ground-state is taken as the initial state, thus giving the same result also for the second ground-state $|\psi_{0}\rangle=|\uparrow\rangle_{z}$.
Other two aspects of the IPR are noteworthy. The first one is the appearance of periodic peaks that decay in amplitude along the dynamics. These peaks are due to partial revivals in the dynamics, that is, the state of the system periodically returns almost completely to the initial state, a feature expected for systems that exhibit DQPT including the LMG model [71, 52]. This fact can also be seen by noticing that the first term of the sum in the IPR is the squared survival probability
$\text{IPR}(t)=\mathcal{L}^{2}(t)+\sum_{k>0}p_{k}^{2}(t),$ (18)
which characterizes DQPT as described in the section II, and the relation above holds for both the pre-quench energy basis and the Krylov basis in our protocol. Therefore, the IPR is also controlled by the survival probability. The second aspect of the IPR that we highlight is the decay rate. We observe that the closer the quench from the dynamical critical point, the more states participate in the dynamics, and hence, the faster the IPR decreases. An increase in the number of Krylov states effectively participating in the dynamics was observed for quenches crossing the equilibrium QPT of the LMG model in the Ref. [32] as well.
The reason for the match when $h_{0}=0$ will be analytically explored later in this article. However, before presenting this analysis, let us consider another interesting quantity, the Shannon entropy.
@@ V.3 $\mathcal{K}$-Entropy - Shannon entropy in the Krylov basis
is a measure of the uncertainty of the system or the classical information contained in the system. We are here interested in the entropy associated with both bases discussed in the previous section: the Krylov basis, for which $p_{n}(t)=|\langle K_{n}|\psi(t)\rangle|^{2}$ and the initial energy basis, whose probabilities are $p_{n}(t)=|\langle E^{0}_{n}|\psi(t)\rangle|^{2}$. The Shannon entropy in the Krylov basis was introduced in Ref. [12] and is also known as the $\mathcal{K}$-entropy. It is a measure of the complexity of the dynamics. The authors of Ref. [35] argued that the complexity defined as the exponential of this quantity measures the minimum Hilbert space dimension required to store the probability distribution $p_{n}(t)$.
By comparing the Shannon entropy in both bases, a perfect match is again observed. Also, such a feature only occurs when we start at $h_{0}=0$. An example is shown in Fig. 4 for the same parameters used in Fig. 3 and we show instances of deviation from this behavior when $h_{0}\neq 0$ in the appendix A.
@@
We note that the probabilities $p_{n}(t)=|\langle E^{0}_{n}|\psi(t)\rangle|^{2}$ are the populations of the density matrix in the energy eigenbasis and, thus, the entropy associated with such distribution is the diagonal entropy, proposed as the thermodynamic entropy for closed quantum systems [72]. Since $p_{n}(t)=|\langle K_{n}|\psi(t)\rangle|^{2}$ can be seen as the probability density associated with site $n$ in the semi-infinite Krylov lattice, naturally, the corresponding Shannon entropy $\mathcal{E}_{\mathcal{K}}(t)$ can be interpreted as the uncertainty in the spreading of the initial state through the Krylov subspace.
The numerical results described above indicate a deeper match beyond the average spread. The goal of the next subsection is to investigate why this happens.
V.4 Derivation of the Krylov basis for $h_{0}=0$
@@
$H_{f}|j,m_{z}\rangle=c_{0}|j,m_{z}\rangle+c_{+}|j,m_{z}+1\rangle+c_{-}|j,m_{z}-1\rangle,$ (23)
@@
The above discussion proves that the Krylov complexity must have the same time behavior as the magnetization. We thus conclude that its time average is an order parameter for this model.
Finally, we note that when $|\psi_{0}\rangle=|j,m_{z}\rangle$ for any $m_{z}$, $\dim{(\mathcal{K})}=\dim{(\mathcal{H})}$, i.e., the dimension of the Krylov subspace is equal to the dimension of the Hilbert space. These initial states explore the whole Hilbert space, showing a kind of ergodicity. We emphasize that this is not a feature of all quantum systems.
@@ VI Conclusion
Understanding the complex nature of the temporal evolution of many-body quantum systems holds fundamental significance across various research fields. This study contributes to this area by examining the Krylov complexity (spread complexity) and its connection to the dynamical phase transition within the LMG model. In essence, we show that the time-averaged Krylov complexity acts as an order parameter in this context. This conclusion stems from a numerical investigation encompassing not only the Krylov complexity but also the inverse participation ratio and Shannon entropy within the Krylov basis. This analysis suggests a relation between the Krylov basis and the energy eigenbasis. A further analytical study establishes that the time-averaged Krylov complexity effectively signals the dynamical phase transition in this model using the equivalence between the Krylov and the energy bases.
It is important to note that a thorough numerical investigation reveals that this connection is valid exclusively in instances where the symmetry of the model is broken, as shown analytcially. This prompts questions about how the Krylov complexity behaves under general changes of symmetry. Addressing this question will certainly deepen our comprehension of the dynamics inherent in many-body systems.
Our findings also provide insights into the thermodynamics of critical systems of this nature. While the thermodynamics of equilibrium phase transitions is well-established, the same cannot be said for its dynamic counterpart. Our results indicate that, in the case of the LMG model, the $\mathcal{K}$-entropy serves as the quantum thermodynamic entropy. This assertion is grounded in the definition of diagonal entropy [72] and the gauge theory of quantum thermodynamics outlined in Ref. [73]. Furthermore, this observation aligns with earlier investigations into the thermodynamics of DQPTs, where a thermodynamic entropy in quantum phase space not only signals the transition but also, on average, exhibits a monotonic increase over time [74]. Consequently, our results suggest a profound connection between DPQTs and the process of thermalization.
@@
Whether the time-averaged Krylov complexity can function as an order parameter for DPT-I in systems beyond the LMG model remains an open question. Additionally, elucidating the broader applicability of the connection between Krylov and energy bases is crucial. Addressing these questions should provide valuable insights into the underlying nature of this dynamical critical behavior.